OpenStudy (anonymous):

HELP ME PLEASE::Jessica is baking a cake. The recipe says that she has to mix 96 grams of sugar to the flour. Jessica knows that 1 cup of this particular sugar has a mass of 128 grams. She added one over two of a cup of sugar to the flour. Should Jessica add more sugar to make the exact recipe or did she go over and by what amount? She went over by two over three of a cup She needs to add two over three of a cup She went over by one over four of a cup She needs to add one over four of a cup

OpenStudy (anonymous):

well if she only added 1/2 cups, that means she only added (128/2) grams of sugar, which is 64 grams. From here you can see that she still needs more sugar so 1 and 3 are already out of the question. Now to find out how much more she needs to add, it's 32 grams. so 32/128= 1/4 and that should give you your answer ;D

OpenStudy (anonymous):

**96g-64g=32 grams**